# Enhanced entanglement scaling and area-law charge fluctuations in a non-Fermi liquid of composite fermions

### v1

@@ Abstract
Discussion.—We have evaluated the scaling of $S_{2}$ Rényi entropy of CFL states at different fillings $\nu{=}1/m$ and showed that it obeys the Widom scaling as in ordinary Fermi liquids but with an enhanced prefactor. We have found an enhancement in both the sphere and torus geometry; moreover, for $m{\geq}2$, the enhancement is essentially unaffected by the LLL projection. We corroborated these conclusions by demonstrating that the wave function [Eq. (1)] encodes the expected area law for charge fluctuations with a universal corner contribution.
Our results for $m{=}2$ are in agreement with Ref. [27], while the discrepancy with Ref. [28] is, at least partly, due to the LLL projection which was not explicitly enforced on the lattice. Moreover, we pointed out the difficulties of interpreting the entropy scalings of the modulus and sign parts of the wave function, arguing that the total entropy is a more robust quantity in the range of numerically accessible system sizes.
@@
One remaining enigma is how to reconcile our results with the iDMRG study of the $\nu{=}1/2$ CFL state on an infinite cylinder, where no enhancement to the Widom formula was found [25]. Two important differences in the setup of Ref. [25] are the quasi-1D geometry, which can impact the behavior of the 2D structure factor [61], and the study of the ground state of the Coulomb interaction rather than a variational wave function. Properties of the CFL for long-range interactions are known to be more similar to those of an ordinary Fermi liquid, while the differences are amplified by shorter-range interactions [2]. While our wave function [Eq. (1)] is not an exact ground state of any known Hamiltonian, it would be interesting to repeat the analysis of Ref. [25] and check the impact of short-range interaction on $\kappa$. Ultimately, the enhanced entanglement scaling in CFL states should be understood analytically. We expect that further progress could be made using multi-dimensional bosonization [62, 63], or by finding other observables whose fluctuations violate the area law, as recently achieved for the gapless Mott insulator with a spinon Fermi surface [47, 31].
@@ Acknowledgements.
Supplemental Online Material for “Enhanced entanglement scaling and area-law charge fluctuationn a non-Fermi liquid of composite fermions”
This Supplemental Material contains (i) expressions for CFL wave functions on the sphere and torus; (ii) an overview of the Jain-Kamilla projection scheme, showing that our results are insensitive to the details of the projection scheme; (iii) analysis of the entanglement scaling and demonstration that the Widom violation is independent of the shape of the subregion; (iv) derivation of the free-fermion Widom formula on the sphere; (v) brief introduction to the SWAP algorithm and our implementation; (vi) analysis of the second Rényi entropy decomposition and charge fluctuations, along with the extraction of their scaling and corner contributions, respectively; (vii) entanglement entropy scaling for flux-attached CFL wave functions.
SI Jain-Kamilla projection for the CFL wave functions on the sphere and torus
In the main text, we have written down the CFL wave function in the infinite disk geometry. For numerical studies of the bulk properties of this wave function, it is more convenient to use compact geometries such as the sphere or torus. Here we explain how the wave functions from the main text can be adapted to such geometries. Since these wave functions do not fully reside in the lowest Landau level, we also briefly describe how the Jain-Kamilla (JK) projection is performed in the two geometries.
SI.1 Torus
On the torus, the single-particle orbitals are plane waves, $\chi_{n}(\mathbf{r}_{j})=\exp(i\mathbf{k}_{n}\cdot\mathbf{r}_{j})$, with $\mathbf{k}_{n}$ the wave vectors of the occupied states of the Fermi sea. The latter can be determined from an empirical rule given in Ref. [38, 37, 27]. The Laughlin wave function $\psi^{\mathrm{L}}_{m}$ is expressed in terms of Jacobi $\vartheta$ functions that depend on complex electron coordinates $z_{j}=x_{j}+iy_{j}$ and modular parameter $\tau$ of the torus [64]:
$\psi^{\mathrm{L}}_{m}=\vartheta\begin{bmatrix}\frac{N-1}{2}\\ \frac{m(N-1)}{2}\end{bmatrix}\bigg{(}\frac{m\sum z_{i}}{L_{1}}\bigg{|}m\tau\bigg{)}\prod_{i<j}\bigg{\{}\vartheta\begin{bmatrix}\frac{1}{2}\\ \frac{1}{2}\end{bmatrix}\bigg{(}\frac{z_{i}-z_{j}}{L_{1}}\bigg{|}\tau\bigg{)}\bigg{\}}^{m}\exp\bigg{(}{\sum_{i}\frac{z_{i}^{2}-|z_{i}^{2}|}{4\ell^{2}}}\bigg{)},\;\;\;\;\;$
with $\ell$ being the magnetic length. We have implicitly chosen the zero momentum sector for the center-of-mass, and strictly periodic conditions for the torus [65]. Here, JK projection can be implemented according to Ref. [66], leading to the following family of projected wave functions:
$\Psi^{\mathrm{CFL,proj[\bm{\alpha}]}}_{m}$ $=\vartheta\begin{bmatrix}\frac{N-1}{2}\\ \frac{m(N-1)}{2}\end{bmatrix}\bigg{(}\frac{m\left(\sum z_{i}+i\ell^{2}\sum k_{i}\right)}{L_{1}}\bigg{|}m\tau\bigg{)}\mathrm{Det}[g_{nl}^{[\bm{\alpha}]}]\exp\bigg{(}{\sum_{i}\frac{z_{i}^{2}-|z_{i}^{2}|}{4\ell^{2}}}\bigg{)}\,,$
$g_{nl}^{[\bm{\alpha}]}$ $=e^{k_{n}(k_{n}+2\overline{k}_{n})\ell^{2}/4}e^{i(k_{n}+\overline{k}_{n})z_{l}/2}\prod_{p=1}^{m/2}\prod_{j\neq l}\vartheta\begin{bmatrix}\frac{1}{2}\\ \frac{1}{2}\end{bmatrix}\bigg{(}\frac{z_{l}+i\alpha_{p}k_{n}\ell^{2}-z_{j}}{L_{1}}\bigg{|}\tau\bigg{)}\,,$
where $g_{nl}^{\bm{\alpha}}$ are CF “single-particle” wave functions that now depend on all other coordinates and $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{m/2})$ are the JK projection coefficients. Torus periodicity enforces the constraint $\sum_{p=1}^{m/2}\alpha_{p}=m$. Note that, essentially, the LLL projection shifts the attached vortices by an amount proportional to the wavevector of the CF. This scheme encompasses all fermionic wave functions with even $m$; for bosons, one needs to attach one additional set of Jastrows outside the wave function for $m\geq 3$ and remove one for $m=1$.
Figure S1: The $S_{2}$ Rényi entropy (left panel) and its decomposition (right panel) for the two choices of JK projection coefficients, $\bm{\alpha}=(2,2)$ and $\bm{\alpha}=(4,0)$, at filling factor $\nu=1/4$. The differences are seen to be minimal, the slope in particular agrees within the error bars, suggesting that details of the JK projection scheme are unimportant in determining the entanglement scaling. All data is for the system size of $N=37$ electrons on the torus.
To verify the consistency of our results for different choices of $\bm{\alpha}$, Fig. S1 shows the $S_{2}$ Rényi entropy scaling is nearly unchanged for two choices $\bm{\alpha}=(2,2)$ and $\bm{\alpha}=(4,0)$ at $\nu=1/4$. Moreover, even the individual contributions to the entropy, $S_{2,p}$, $S_{2,\mathrm{mod}}$ and $S_{2,\mathrm{sign}}$, discussed below, are essentially the same for the two $\bm{\alpha}$ choices. This shows that $S_{2}$ Rényi entropy is insensitive to the details of JK projection coefficients, unlike some other quantities such as the Hall viscosity in Ref. [67].
SI.2 Sphere
On the sphere, the single-particle orbitals are the standard spherical harmonics, $\chi_{n}(\mathbf{r}_{j})=Y_{L_{n}}^{M_{n}}(\theta_{j},\varphi_{j})$, where $L$, $M$ stand for the angular momentum and its $z$-component, respectively, while $\theta_{j}$, $\phi_{j}$ are the standard spherical angles describing the position of $j$th particle. The Laughlin wave function,
$\psi^{\mathrm{L}}_{m}=\prod_{i<j}(u_{i}v_{j}-v_{i}u_{j})^{m},$ (S1)
is expressed in terms of spherical spinor coordinates, $u_{j}{=}\sin(\theta_{j}/2)\exp(i\phi_{j}/2)$ and $v_{j}{=}\cos(\theta_{j}/2)\exp(-i\phi_{j}/2)$ [68]. Thus, the unprojected CFL wave function on the sphere has a very similar form to that written in the main text for the infinite plane. Fig. S2 demonstrates the similarity of unprojected CFL results on the sphere with those presented in the main text for the torus geometry. This consistency shows the robustness of our results, e.g., for the sequence of system sizes $N=n^{2}$ we consider on the sphere, the CFL has a uniform $L{=}0$ state, which removes some of the ambiguity associated with the definition of the Fermi momentum on the torus.
Figure S2: Second Rényi entropy of unprojected CFL states on the sphere at fillings $\nu{=}1/m$, with $m{=}1,2,3,4$. All data is for system size $N{=}64$. Crosses represent Monte Carlo data, while the continuous line for free fermions represents ED data. All filling factors show an enhanced slope compared to the Widom value.
Finally, we explain how to perform the projection on the sphere. Let us first treat the state at $\nu{=}1/2$, with two vortices attached i.e., $\Psi^{\rm CFL}_{1/2}{=}\mathcal{P}_{\rm LLL}\prod_{j{<}k}(u_{j}v_{k}{-}u_{k}v_{j})^{2}\Phi_{\rm FS}$, where $\Phi_{\rm FS}$ is the filled-shell wave function on the sphere at zero-field that is analogous to that of a Fermi sea [26]. The projected state can be evaluated as a Slater determinant of CF wave functions, which now depend on the coordinates of all particles through the Jastrow factors $J_{j}{=}\prod_{k{\neq}j}(u_{j}v_{k}{-}u_{k}v_{j})$. We quote the final expression here:
$Y^{CF}_{0LM}(\mathbf{\Omega}_{j})=N_{0,L,M}(-1)^{L-M}\frac{(2Q+1)!}{(2Q+L+1)!}u_{j}^{M}v_{j}^{-M}\sum_{s=0}^{L}(-1)^{s}\begin{pmatrix}L\\ s\end{pmatrix}\begin{pmatrix}L\\ L-M-s\end{pmatrix}u_{j}^{s}v_{j}^{L-s}\bigg{[}\bigg{(}\frac{\partial}{\partial u_{j}}\bigg{)}^{s}\bigg{(}\frac{\partial}{\partial v_{j}}\bigg{)}^{L-s}J_{j}\bigg{]}\,,$ (S2)
where $N_{0,L,M}$ is a normalization factor, and the flux for the physical electrons is $2Q{=}2(N{-}1)$. Given that we only have one power of Jastrow, the derivatives can be evaluated as follows:
$\bigg{(}\frac{\partial}{\partial u_{j}}\bigg{)}^{s}\bigg{(}\frac{\partial}{\partial v_{j}}\bigg{)}^{L-s}J_{j}=J_{j}\sum\prod_{i=0}^{s}\frac{v_{k_{i}}}{u_{j}v_{k_{i}}-u_{k_{i}}v_{j}}\prod_{i=0}^{L-s}\frac{-u_{l_{i}}}{u_{j}v_{l_{i}}-u_{l_{i}}v_{j}}\,,$ (S3)
where the sum runs over all possibilities in choosing $s$ particles $\{k_{i}\}$ and $L{-}s$ particles $\{l_{i}\}$, all distinct. The complete Jastrow factor $J_{i}$ can be factored out of the determinant post-projection.
Note that the previous wave function directly allows construction for CFL states at different fillings. For the fermionic state at $\nu{=}1/4$, we choose to only project with two Jastrow factors, leaving the other two outside – this is similar to the $\bm{\alpha}=(4,0)$ state on the torus. For the bosonic states at $\nu{=}1$ and $\nu{=}1/3$, we need to divide or multiply by a single Jastrow factor respectively [69, 70], again mirroring the construction on the torus. It is known that the microscopic CFL wave function is not very sensitive to the number of Jastrows placed inside the LLL-projection operator [71, 72].
SII Slope extrapolation and shape dependence of the entanglement scaling
Here we present additional data on the slope extrapolations used in Fig. 1(d) of the main text. Fig. S3(a) shows the data points used at system sizes $N\in\{12,21,32,37,60\}$ for the projected CFL wave function at $\nu{=}1/2$. We fix the smallest value to $\ln\lambda\approx 0.5$ regardless of system size, while the end point increases with $N$ such that $\ln\lambda_{\mathrm{end}}\in\{1.33,1.51,1.59,1.66,1.75\}$. These values are chosen to maximize the range of data points used at each system size, while still correctly extrapolating the correct Widom coefficient in the thermodynamic limit (see inset).
Fig. S3(b) and (c) show a comparison between systems with subregions of circular and square shape, respectively. We find the ratio of the slopes to be approximately independent of the subregion geometry. This indicates that the CFL entanglement scaling still encodes the geometric information in the Widom formula, with the violation limited to a numerical overall prefactor.
Figure S3: Slope extrapolation and shape dependence of the entanglement scaling. (a): We extract the $S_{2}$ slope from a subset of data points in the $\lambda$-range shown here for various system sizes indicated in the legend. Solid lines are free-fermion data, while markers denote the projected CFL data at $\nu{=}1/2$. The range of $\lambda$ values is chosen with free fermions as a benchmark, such that the correct Widom slope is recovered in the thermodynamic limit (inset). (b)-(c): The dependence of the entanglement scaling on the shape of the subregion, i.e., circle in (b) or square in (c). $S_{2}$ for free fermions (solid lines) and the projected CFL state at $\nu=1/2$ (markers), for system size $N{=}37$. The ratio of slopes is the approximately independent of subregion geometry.
SIII Free fermions
If a wave function $\Psi$ can be written as a Slater determinant, $\Psi={\rm Det}[\phi_{m}(\mbox{\boldmath$r$}_{n})]$, where the single-particle orbitals $\phi_{m}$ are orthonormalized over the full space, one can efficiently compute the entanglement entropy using the correlation matrix [73]. The correlation matrix $\mathcal{A}$ has matrix elements given by the overlap integrals on the subsystem, $\mathcal{A}_{mn}=\int_{A}\mathrm{d}\mathbf{r}\,\phi_{m}^{*}(\mathbf{r})\phi_{n}(\mathbf{r})$. Suppose the eigenvalues of $\mathcal{A}$ are $a_{m}$, $m=1,2,\cdots N$; the 2nd Renyi entropy is given by [74, 75]
$S_{2}=-\sum_{m}\ln\left(a_{m}^{2}+(1-a_{m})^{2}\right).$ (S4)
@@
In the torus geometry with a circular subregion $A$, the overlap matrix elements evaluate to:
$\mathcal{A}_{mn}^{\mathrm{torus}}=\int_{A}\mathrm{d}\mathbf{r}\,\phi_{m}^{*}(\mathbf{r})\phi_{n}(\mathbf{r})=\frac{1}{L^{2}}\int_{A}\mathrm{d}\mathbf{r}\,e^{i(\mathbf{k}_{n}-\mathbf{k}_{m})\cdot\mathbf{r}}=\frac{2\pi}{L^{2}}\frac{r_{A}}{|\mathbf{k}_{n}-\mathbf{k}_{m}|}J_{1}(|\mathbf{k}_{n}-\mathbf{k}_{m}|r_{A})\,,$ (S5)
@@
where $J_{1}$ is the Bessel function of the first kind. On the sphere, we obtain for a spherical cap subregion
$\mathcal{A}_{mn}^{\mathrm{sphere}}=\int_{A}\mathrm{d}\Omega\,Y^{*}_{L_{m}M_{m}}(\Omega)Y_{L_{n}M_{n}}(\Omega)=2\pi\mathcal{N}_{L_{m}M_{m}}\mathcal{N}_{L_{n}M_{n}}\delta_{M_{m},M_{n}}\int_{\cos\theta_{A}}^{1}\mathrm{d}x\,P_{L_{m}}^{M_{m}}(x)P_{L_{n}}^{M_{n}}(x)\,,$ (S6)
@@
SIV The Widom formula on the sphere

### v2

@@ V Conclusions and discussion
We have evaluated the scaling of $S_{2}$ Rényi entropy of CFL states at different fillings $\nu{=}1/m$ and showed that it obeys the Widom scaling as in ordinary Fermi liquids but with an enhanced prefactor. We have found an enhancement in both the sphere and torus geometry; moreover, for $m{\geq}2$, the enhancement is essentially unaffected by the LLL projection. We have corroborated these conclusions by demonstrating that the wave functions in Eq. (1) encode the expected area law for charge fluctuations with a universal corner contribution. Moreover, we pointed out the difficulties of interpreting the entropy scalings of the modulus and sign parts of the wave function, arguing that the total entropy is a more robust quantity in the range of numerically accessible system sizes.
Our results for $m{=}2$ are in agreement with Ref. [27], while the discrepancy with Ref. [28] is, at least partly, due to the LLL projection which was not explicitly enforced on the lattice. In fact, our unprojected results in Fig. 5(b) are qualitatively similar to Ref. [28]; by contrast, the LLL-projected results in Fig. 1 show a visible departure from Ref. [28], e.g., the entropy of the bosonic $\nu{=}1$ CFL is similar to that of the fermionic $\nu{=}1/2$ CFL state after LLL projection, unlike their lattice analogs. These differences point to a subtle role of LLL projection, precluding direct comparisons between continuum and lattice versions of the CFL state. This could potentially imply that the entanglement scaling may depend on the microscopic details of flux attachment (see Appendix F). Unfortunately, as we explained, settling this question numerically requires access to system sizes several times larger than the current computational facility.
@@
One remaining enigma is how to reconcile our results with the iDMRG study of the $\nu{=}1/2$ CFL state on an infinite cylinder, where no enhancement to the Widom formula was found [25]. Two important differences in the setup of Ref. [25] are the quasi-1D geometry, which can impact the behavior of the 2D structure factor [77], and the study of the ground state of the Coulomb interaction rather than a variational wave function. Properties of the CFL for long-range interactions are known to be more similar to those of an ordinary Fermi liquid, while the differences are amplified by shorter-range interactions [2]. While our wave function [Eq. (1)] is not an exact ground state of any known Hamiltonian, it would be interesting to repeat the analysis of Ref. [25] and check the impact of short-range interaction on $\kappa$.
Ultimately, the enhanced entanglement scaling in CFL states should be understood analytically. We expect that further progress could be made using multi-dimensional bosonization [78, 79], or by finding other observables whose fluctuations violate the area law, as recently achieved for the gapless Mott insulator with a spinon Fermi surface [61, 32]. This understanding may shed light on other pertinent questions, such as whether the Chern band projection similarly enhances the entanglement of a CFL in zero magnetic field [6, 7], or whether the latter also depends on band geometry and Chern number.
@@ Acknowledgements.
Figure 6: Illustration of the SWAP process. We have two copies of the full system, whose particles are labeled with different colors, and each copy is divided into two regions separated by the dashed circle. Through the SWAP process, we interchange the particle coordinates inside the circle between the two copies while keeping the outside coordinates unchanged.
Appendix A SWAP algorithm and entropy decomposition
Following Refs. [31, 72], in our Monte Carlo evaluations of $S_{2}$ Rényi entropy we replace the partial trace with an expectation value $\langle\mathrm{SWAP}_{A}\rangle$ of the SWAP operator between two copies of the system, as illustrated in Fig. 6. This expectation value can be naturally broken down into a product of three terms:
$\langle\mathrm{SWAP}_{A}\rangle=P_{\mathrm{SWAP}}\,\langle\mathrm{SWAP}_{\mathrm{mod}}\rangle\langle\mathrm{SWAP}_{\mathrm{sign}}\rangle\,.$ (15)
$P_{\mathrm{SWAP}}$ $=\frac{\int\mathrm{d}\bm{\alpha}\mathrm{d}\bm{\beta}\,|\Psi(\alpha_{1},\beta_{1})|^{2}|\Psi(\alpha_{2},\beta_{2})|^{2}\delta_{N_{1},N_{2}}}{\int\mathrm{d}\bm{\alpha}\mathrm{d}\bm{\beta}\,|\Psi(\alpha_{1},\beta_{1})|^{2}|\Psi(\alpha_{2},\beta_{2})|^{2}}\,,$ (16)
The last two terms in Eq. (15) are the “mod” and “sign” contributions, which are evaluated only in the subspace of swappable copies ($N_{1}=N_{2}$),
$\langle\mathrm{SWAP}_{\mathrm{mod}}\rangle$ $=\frac{\int\mathrm{d}\bm{\alpha}\mathrm{d}\bm{\beta}\,\delta_{N_{1},N_{2}}|\psi(\alpha_{1},\beta_{1})|^{2}|\psi(\alpha_{2},\beta_{2})|^{2}\left|\frac{\psi(\alpha_{2},\beta_{1})\psi(\alpha_{1},\beta_{2})}{\psi(\alpha_{1},\beta_{1})\psi(\alpha_{2},\beta_{2})}\right|}{\int\mathrm{d}\bm{\alpha}\mathrm{d}\bm{\beta}\,\delta_{N_{1},N_{2}}|\psi(\alpha_{1},\beta_{1})|^{2}|\psi(\alpha_{2},\beta_{2})|^{2}},$ (17)
$\langle\mathrm{SWAP}_{\mathrm{sign}}\rangle$ $=\frac{\int\mathrm{d}\bm{\alpha}\mathrm{d}\bm{\beta}\,\delta_{N_{1},N_{2}}|\psi(\alpha_{2},\beta_{1})\psi(\alpha_{1},\beta_{2})\psi(\alpha_{1},\beta_{1})\psi(\alpha_{2},\beta_{2})|\,e^{i\theta(\bm{\alpha},\bm{\beta})}}{\int\mathrm{d}\bm{\alpha}\mathrm{d}\bm{\beta}\,\delta_{N_{1},N_{2}}|\psi(\alpha_{2},\beta_{1})\psi(\alpha_{1},\beta_{2})\psi(\alpha_{1},\beta_{1})\psi(\alpha_{2},\beta_{2})|}\;.$ (18)
$\exp i\theta(\bm{\alpha},\bm{\beta})=\frac{\psi(\alpha_{1},\beta_{2})^{*}\psi(\alpha_{2},\beta_{1})^{*}\psi(\alpha_{1},\beta_{1})\psi(\alpha_{2},\beta_{2})}{|\psi(\alpha_{1},\beta_{2})\psi(\alpha_{2},\beta_{1})\psi(\alpha_{1},\beta_{1})\psi(\alpha_{2},\beta_{2})|}\;.$ (19)
$S_{2}=S_{2,\mathrm{p}}+S_{2,\mathrm{mod}}+S_{2,\mathrm{sign}}.$ (20)
In previous works, such as Ref. [28], the first two contributions to the entropy were lumped together, because $S_{2,\mathrm{p}}+S_{2,\mathrm{mod}}\equiv S_{2}(|\Psi|)$ can be interpreted as the entropy of the absolute value of the wave function $\Psi$. While this may be useful in certain cases [73], our data below suggests this is not helpful for gapless CFL states since $S_{2,\mathrm{p}}$ and $S_{2,\mathrm{mod}}$ exhibit different scaling with $\lambda$. Hence, we study individually the three distinct contributions to $S_{2}$.
Appendix B Overlap integrals
@@
To evaluate $S_{2}$ Rényi entropy for free fermions in a finite-size system, we use Eq. (3) which requires knowledge of the overlap integrals on the subsystem, $\mathcal{A}_{mn}$. In the torus geometry with a circular subregion $A$, the overlap matrix elements evaluate to:
$\mathcal{A}_{mn}^{\mathrm{torus}}$ $=\int_{A}\mathrm{d}\mathbf{r}\,\phi_{m}^{*}(\mathbf{r})\phi_{n}(\mathbf{r})=\frac{1}{L^{2}}\int_{A}\mathrm{d}\mathbf{r}\,e^{i(\mathbf{k}_{n}-\mathbf{k}_{m})\cdot\mathbf{r}}$
$=\frac{2\pi}{L^{2}}\frac{r_{A}}{|\mathbf{k}_{n}-\mathbf{k}_{m}|}J_{1}(|\mathbf{k}_{n}-\mathbf{k}_{m}|r_{A})\,,$ (21)
@@
where $J_{1}$ is the Bessel function of the first kind. On the sphere, for a spherical cap subregion, we obtain
$\mathcal{A}_{mn}^{\mathrm{sphere}}$ $=\int_{A}\mathrm{d}\Omega\,Y^{*}_{L_{m}M_{m}}(\Omega)Y_{L_{n}M_{n}}(\Omega)$
$=2\pi\mathcal{N}_{L_{m}M_{m}}\mathcal{N}_{L_{n}M_{n}}\delta_{M_{m},M_{n}}$
×$\times\int_{\cos\theta_{A}}^{1}\mathrm{d}x\,P_{L_{m}}^{M_{m}}(x)P_{L_{n}}^{M_{n}}(x)\,,$ (22)
@@
Appendix C Comparison of Rényi entropies obtained with different Jain-Kamilla projections
To verify the consistency of our results for different choices of $\bm{\alpha}$, Fig. 7 shows the $S_{2}$ Rényi entropy scaling is nearly unchanged for two choices $\bm{\alpha}=(2,2)$ and $\bm{\alpha}=(4,0)$ at $\nu=1/4$. Moreover, even the individual contributions to the entropy, $S_{2,p}$, $S_{2,\mathrm{mod}}$ and $S_{2,\mathrm{sign}}$, discussed in Appendix A, are essentially the same for the two $\bm{\alpha}$ choices. This shows that $S_{2}$ Rényi entropy is insensitive to the details of JK projection coefficients, unlike some other quantities such as the Hall viscosity in Ref. [81].
Figure 7: The $S_{2}$ Rényi entropy (left panel) and its decomposition (right panel) for the two choices of JK projection coefficients, $\bm{\alpha}=(2,2)$ and $\bm{\alpha}=(4,0)$, at filling factor $\nu=1/4$. The differences are seen to be minimal, the slope in particular agrees within the error bars, suggesting that details of the JK projection scheme are unimportant in determining the entanglement scaling. All data is for the system size of $N=37$ electrons on the torus.
Appendix D Slope extrapolation and shape dependence of the entanglement scaling
Here we present additional data on the slope extrapolations used in Fig. 1(c) of the main text. Fig. 8(a) shows the data points used at system sizes $N\in\{12,21,32,37,60\}$ for the projected CFL wave function at $\nu{=}1/2$. We fix the smallest value to $\ln\lambda\approx 0.5$ regardless of system size, while the endpoint increases with $N$ such that $\ln\lambda_{\mathrm{end}}\in\{1.33,1.51,1.59,1.66,1.75\}$. These values are chosen to maximize the range of data points used at each system size, while still correctly extrapolating the correct Widom coefficient in the thermodynamic limit – see inset of Fig. 8(a).
Fig. 8(b) and (c) show a comparison between systems with subregions of circular and square shape, respectively. We find the ratio of the slopes to be approximately independent of the subregion geometry. This indicates that the CFL entanglement scaling still encodes the geometric information in the Widom formula, with the violation limited to a numerical overall prefactor.
Figure 8: Slope extrapolation and shape dependence of the entanglement scaling. (a): We extract the $S_{2}$ slope from a subset of data points in the $\lambda$ range shown here for various system sizes indicated in the legend. Solid lines are free-fermion data, while markers denote the projected CFL data at $\nu{=}1/2$. The range of $\lambda$ values is chosen with free fermions as a benchmark, such that the correct Widom slope is recovered in the thermodynamic limit (inset). (b)-(c): The dependence of the entanglement scaling on the shape of the subregion, i.e., circle in (b) or square in (c). $S_{2}$ for free fermions (solid lines) and the projected CFL state at $\nu=1/2$ (markers), for system size $N{=}37$. The ratio of slopes is approximately independent of subregion geometry.
Appendix E The Widom formula on the sphere